OpenStudy (anonymous):

use power reduction formula to rewrite the equation in terms of cosine: cos^6 x

OpenStudy (anonymous):

here u can use some formula like cos 3x = 3cos x - 4cos^3x.........(1) cos^2x = (cos2x +1)/2.............(2) ans: cos^6x= (cos^3x)^2 = ( (3cosx-cos 3x)/4)^2 =(9co^2x+cos^2 (3x) -6 cos3x* cos x)/16 =[9(1+cos2x)/2 + (1+cos 6x)/2 -6cos3x* cos x]/16 =[ 9/32( 1+ cos2x) +1/32( 1+cos 6x) - 3/8(cos3x * cos x)

OpenStudy (nowhereman):

from wikipedia: \[\cos^n\theta = \frac{1}{2^n} \binom{n}{\frac{n}{2}} + \frac{2}{2^n} \sum_{k=0}^{\frac{n}{2}-1} \binom{n}{k} \cos{((n-2k)\theta)} \] because n is even. Insert n=6 and ready you are.

OpenStudy (anonymous):

this u can go for direct.

OpenStudy (anonymous):

Still here, rinsped?